OpenStudy (anonymous):

(2x-1)^2 how do you solve

OpenStudy (anonymous):

You distribute the 2 so it becomes 4x-2

OpenStudy (anonymous):

4x-2

OpenStudy (anonymous):

would you square it?

OpenStudy (anonymous):

actually it is 4x^2-2

OpenStudy (anonymous):

who is Right?

OpenStudy (anonymous):

actually it is 4x^2 - 4x + 1

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

because (2x-1)^2 = (2x-1)(2x-1)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

something squared is that thing times itself so (2x-1)^2 = (2x-1)(2x-1) when you multiply it out you get 4x^2 - 2x -2x +1 =4x^2 - 4x + 1

OpenStudy (anonymous):

ok